OpenStudy (phagey):

When solving a system of equations using elimination, it is necessary to make sure the coefficients in front of one of the variables are additive inverses. A. True B. False

jhonyy9 (jhonyy9):

how you think it - do you can writing here one example please ?

jhonyy9 (jhonyy9):

2x+3y=4 x+5y=9 solve this system with elimination method

OpenStudy (anonymous):

False

jhonyy9 (jhonyy9):

@TheDebunker do you understand the question right,correct ?

OpenStudy (phagey):

its true?

jhonyy9 (jhonyy9):

@phagey how you think it ?

OpenStudy (phagey):

true? @jhonyy9

jhonyy9 (jhonyy9):

are you sure ? why not can being false ?

OpenStudy (phagey):

im really not sure @jhonyy9

jhonyy9 (jhonyy9):

2x+3y=4 x+5y=9 using elimination method you need multiplie or make one variable so that adding them than result zero - in this case multiplie second equation by -2 so will be eliminated the terms with x - yes ?

jhonyy9 (jhonyy9):

2x+3y =4 -2x -10y=-18 ------------- 0 -7y = -14 right ? so how you think it now after you have got this example ?

OpenStudy (phagey):

true

jhonyy9 (jhonyy9):

2x and -2x how are them not additive inverses ?

jhonyy9 (jhonyy9):

yes but than you know anything that is right so you need to be sure indifferent what anybody write wrong answer - ok ,? good luck bye bye

OpenStudy (phagey):

thanks

jhonyy9 (jhonyy9):

np welcome was my pleasure bye bye